OpenStudy (nicole143):

@mathmale Just number 13. I only got as far as writing down the question and putting 36 / 9 = 4.

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

9/36 = 1/4

OpenStudy (mathmale):

OK: Let's focus on #13. It's true that 36/9=4, but under the radical, we have 9/36=1/4. \[\sqrt{\frac{ 9x ^{4} }{ 36 }}=\frac{ \sqrt{9} \sqrt{x ^{4}}}{ \sqrt{36} }=?\]

OpenStudy (anonymous):